Calculate the percent of BaO in 29 gram of a mixture BaO and CaO ?Calculate the percent of BaO in 29 gram of a mixture BaO and CaO which just reacts with 100.8 mL of 6.00 M (molarity) HCl ? Rxn is BaO + 2 HCl gives= BaCl2 + H2O and CaO + 2HCl gives= CaCl2 +H2O Kindly solve using stoichiometry and explain

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
We have,
(0.1008 L) x (6.00 mol/L HCl) x (1 mol oxide / 2 mol HCl) = 0.3024 mol oxides.

From 100% BaO, 29 g would contain 
(29 g) / (153.3271 g BaO/mol) = 0.1891 mol oxide .

If the mixture was 100% CaO (0% BaO), 29 g would contain: 
(29 g) / (56.0778 g CaO/mol) = 0.5171 mol oxide 

From unitary method:
If 

0.1891 mol = 100% BaO 
0.3024 mol = z% BaO 
0.5171 mol = 0% BaO 
then
z / 100 = (0.5171 - 0.3024) / (0.5171 - 0.1891) 
z = 65.4% BaO.
This would be the proprtion of BaO in actual.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)